# Learning Policies for Dynamic Coalition Formation in Multi-Robot Task Allocation

### v1

@@ Abstract
We propose a decentralized, learning-based framework for dynamic coalition formation in Multi-Robot Task Allocation (MRTA). Our approach extends Multi-Agent Proximal Policy Optimization (MAPPO) by incorporating spatial action maps, robot motion control, task allocation revision, and intention sharing to enable effective coalition formation. Extensive simulations demonstrate that our model significantly outperforms existing methods, including a market-based baseline. Furthermore, we assess the scalability and generalizability of the proposed framework, highlighting its ability to handle large robot populations and adapt to diverse task allocation environments.
@@ I Introduction
Coalition formation, widely observed in nature, such as in ant colonies and bee hives, inspires researchers and engineers to design simple robots capable of strategic cooperation. In multi-robot systems, coalitions emerge when the collective actions of robots exceed what they could accomplish individually, as illustrated in Fig. 1. In this work, we focus on developing a team of robots capable of performing tasks that require coalition in dynamic environments. In such scenarios, robots must learn to coordinate by leveraging both local information gathered from their own sensors and shared data from neighboring robots to effectively execute complex multi-robot tasks.
Multi-Robot Task Allocation (MRTA) is a research theme directly relevant to our work, which seeks to address the challenge of assigning and executing multiple tasks across a team of robots to optimize a collective objective. We focus on MRTA problems characterized by: (i) each robot being capable of executing one task at a time, (ii) tasks requiring multiple robots simultaneously to be completed, and (iii) robots needing to be adjacent to a task to execute it, requiring movement to the task location beforehand. This problem formulation falls under the Single-Task robots, Multi-Robot tasks, Time-extended Assignment (ST-MR-TA) category of MRTA, as defined by the taxonomy in [1]. Relevant applications include transportation [2] and service dispatch scenarios [3, 4] such as firefighting and large-scale disaster response.
Figure 1: The firetruck dispatch is an illustrative example of a dynamic coalition formation problem: without a centralized coordinator, a team of firetrucks must form coalitions in a decentralized fashion to tackle tasks, such as extinguishing fires, that cannot be accomplished individually.
@@
We propose a learning-based framework for multi-robot dynamic coalition formation to address problems within the ST-MR-TA MRTA category. In this framework, each robot utilizes a receding horizon-type planning strategy to generate a long-term plan for task execution, which is repeatedly revised as necessary in subsequent time steps. The robot shares the plan with other robots that are within its communication range. This combination of task reassessment and sharing of the task execution plan enables dynamic coalition formation. Our contributions are summarized as follows:
@@
    Each robot’s policy is modeled as an end-to-end convolutional neural network based on the U-Net architecture [9]. The network processes task information within the robot’s sensing range and incorporates the intentions of neighboring robots. This shared network is trained using the MAPPO algorithm and then distributed across all robots.
@@
    Through extensive simulations, we demonstrate that our approach outperforms existing methods including the market-based approach, PCFA. Additionally, we demonstrate the scalability of models up to 1000 robots and their generalizability across a wide range of task settings.
@@ II Related Work
Solutions to ST-MR-TA MRTA include optimization-based [10, 11, 12, 13], trait-based [14, 15], and market-based [6, 8, 16, 7] methods. Among decentralized solutions for MRTA, market-based approaches are particularly popular. Specifically, the PCFA [8] is a market-based solution designed for coalition formation in homogeneous teams with dynamic tasks. Due to the similarity in problem formulation, we selected PCFA as the baseline for our simulations reported in §V.
Reinforcement Learning (RL) focuses on learning a policy (i.e., a mapping from observations to actions) that maximizes a robot’s total reward in a Markov decision process. It has been widely applied to various robotics problems, particularly in object manipulation [17, 18, 19, 20, 21]. While many approaches aim to directly control robot actuators, studies such as [17, 18, 22] propose the use of an abstract action map. In this framework, the policy selects a task at each step, while a motion planner handles the low-level actuator control to navigate the robot to the task location. This abstraction alleviates the learned policy from handling low-level control, allowing it to concentrate on long-term planning.
@@
Several studies have applied MARL to MRTA [29, 34, 35, 23], but only [29, 35] have addressed MRTA problems involving multi-robot tasks. The MRTA problem was introduced to the MARL community as a benchmark in [29], where the authors released a Python package called Level-Based Foraging (LBF) [29]. In LBF, each object is assigned a level, and it can only be picked up by a group of robots if the sum of the robots’ levels meets or exceeds the object’s level. Although LBF has become a popular MARL benchmark [36, 37, 38], it lacks support for dynamic object spawning and abstract action spaces (e.g., selecting task positions on the grid). In [36], several MARL algorithms were evaluated across popular benchmarks, including LBF (in fully observable scenarios only), and the EPyMARL package was released for community use. We expand their implementation by incorporating robot communication and additional features, which are detailed further in §IV.
@@
Figure 2: Overview of the robot-environment interaction cycle. On the right a depiction of the environment is shown, along with the state representation used by the critic during training. An observation is given to each robot, where each channel encodes a different type of occupation. The observation is processed by the policy $\pi_{\phi}$ and filtered such that only visible tasks are considered. A task is selected from the distribution and the motion planner finds a feasible path to it. The first step along the path is fed to the environment, and the cycle repeats.
@@ III Problem Description
Each task is associated with a positive integer $l_{i}\in\{1,\dots,L_{max}\}$, referred to as the task’s level, which represents the number of robots required to execute it. Here, $L_{max}$ denotes the maximum task level. We consider a dynamic task set: tasks are removed from the environment once completed by robots, and new tasks can appear at any time step in unoccupied cells. At any time step $t$, the set of available tasks is represented as $\mathbb{T}_{t}=\{(q_{i},l_{i})\}_{i=1}^{T_{t}}$, where $q_{i}$ denotes the location of the $i$-th task, $l_{i}$ indicates its level, and $T_{t}$ represents the total number of tasks available at time $t$.
@@
To design a decentralized policy under partial observability, we model the problem as a Decentralized Partially-Observable Markov Decision Process (Dec-POMDP) [39, 40]. A Dec-POMDP is defined by the tuple $(N,\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},O,R)$, where $N$ is the number of robots, $\mathcal{S}$ represents the state space of the environment, and $\mathcal{A}$ and $\mathcal{O}$ denote the action and observation spaces of individual robots, respectively. The state transition function $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}_{+}$ defines the probability $\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$ of transitioning to the next state $s_{t+1}\in\mathcal{S}$, given the current state $s_{t}\in\mathcal{S}$ and the joint action $\mathbf{a}_{t}\in\mathcal{A}^{N}$ of all robots. The observation function $O:\mathcal{S}\times\{1,\dots,N\}\to\mathcal{O}$ provides each robot $i$ with an observation $o_{i,t}=O(s_{t},i)$, mapping the state space $\mathcal{S}$ and robot index $i$ to the observation space $\mathcal{O}$. Lastly, the shared reward function is represented by $R:\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}$, where $r_{t}=R(s_{t},\mathbf{a}_{t})$.
@@ IV Task Allocation Policy for Dynamic Coalition Formation
The main contributions of this work are twofold: the development of decentralized task allocation policy models and the design of a learning-based algorithm within the Dec-POMDP framework to identify the optimal policy for the proposed models. Both aspects are elaborated in detail in this section. Our proposed policy model is designed to select tasks for execution, assess whether a robot’s current task allocation requires revision, and determine new task assignments based on the robot’s evolving local information. Once a task is assigned, each robot executes low-level control inputs to move along a collision-free path to a neighboring cell of its assigned task within a single time step. In the subsequent time step, the policy reassesses the task allocation based on updated local information, redirecting the robot to a new task if necessary. This iterative process ensures that the overall collective objective is optimized.
IV-A Multi-Agent Proximal Policy Optimization (MAPPO)
To compute the optimal policy, we adopt MAPPO [27], a CTDE algorithm designed for MARL. MAPPO enables robots to share experiences during the learning phase, which is particularly beneficial for training policies in teams of homogeneous robots. MAPPO is an Actor-Critic algorithm comprising two distinct networks: the actor $\pi_{\phi}$ and the critic $V_{\theta}$. The actor outputs a distribution over actions based on each robot’s observations, formally defined by the policy $\pi_{\phi}:\mathcal{O}\times\mathcal{A}\to\mathbb{R}_{+}$, where $\phi$ represents the parameters of the policy network. These parameters $\phi$ are shared across all robots, a technique known as parameter sharing [41], which improves training efficiency, especially in cooperative scenarios involving homogeneous robots. The critic computes an approximate value function $V_{\theta}:\mathcal{S}\to\mathbb{R}$, parameterized by $\theta$, to estimate the expected discounted sum of future rewards. This is expressed as $\mathbb{E}\big{[}{\sum_{\tau=t}^{T}\gamma^{\tau-t}R(s_{\tau},\mathbf{a}_{\tau})}\big{]}$.
The critic $V_{\theta}$ is trained to accurately estimate future rewards, while the actor $\pi_{\phi}$ is trained to select actions that maximize the critic’s estimate. In our framework, both the actor $\pi_{\phi}$ and the critic $V_{\theta}$ are implemented using U-Nets [9], a Convolutional Neural Network (CNN)-based architecture. For this implementation, the actor’s input (observations) and output (action distribution) as well as the critic’s input (state), are represented as images (3D arrays). The U-Net architecture is particularly advantageous due to its ability to capture both local spatial relationships and global context within the input image, making it well-suited for policy learning.
IV-B Dec-POMDP and Task Allocation Policy Model
IV-B1 Reward Function
@@
Higher-level tasks, which require coalitions of multiple robots, are more challenging to complete. To prevent robots from favoring lower-level tasks, we incentivize higher-level tasks by rewarding robots with a value equal to the square of the task level. Specifically, the reward is defined as $r_{t}=\sum_{i\in\mathbb{C}_{t}}l_{i}^{2}$, where $\mathbb{C}_{t}=\mathbb{T}_{t}\backslash\mathbb{T}_{t+1}$ represents the set of tasks completed at time step $t$, and $l_{i}$ denotes the level of task $i$.
IV-B2 State and Observation Spaces
@@
We adopt a 2D occupancy grid representation segmented into multiple channels (see Table I), enabling the use of CNN architectures for policy design, similar to approaches in [34, 17, 23]. This differs from the representation used in LBF [29], which encodes the state as a vector of robots and their positions. In our method, the state is defined as multiple separate channels as $s_{t}\triangleq(\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t},\mathcal{C}^{\text{intent}}_{t})$, where $\mathcal{C}^{\text{intent}}_{t}=\mathcal{C}^{\text{intent}}_{1,t}\times\cdots\times\mathcal{C}^{\text{intent}}_{N,t}$. [^1] To simplify the design and enhance parameter efficiency in training the critic model, the input channel $\mathcal{C}^{\text{intent}}_{t}$ is implemented as a single aggregated channel that combines the intention channels of all robots.
@@
The observation $O(s_{t},i)$ for robot $i$ is derived by projecting and cropping $s_{t}$ onto areas centered around the robot’s position, determined by its view range $R_{\text{view}}$ and communication range $R_{\text{comm}}$. Specifically, the robot intention channel $\mathcal{C}^{\text{intent}}_{t}$ includes $\mathcal{C}^{\text{intent}}_{j,t}$ from robot $j$ if $j$ is within $R_{\text{comm}}$ of robot $i$. Meanwhile, the other channels, $\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t}$, are cropped based on $R_{\text{view}}$.
TABLE I: Description of CNN channels.
@@
IV-B3 Action Space, Motion Planner, and Intention Map
Following a similar representation as in [17, 18], a robot’s action – determined by the task allocation policy $\pi_{\phi}$ – is defined as the location of its next task. Robots can select task locations only if they are within its view range; otherwise all grid locations are available for task location selection. This constraint provides prior information, potentially reducing training time while encouraging robots to commit to specific tasks. By sharing task allocation information with neighboring robots, this approach facilitates coalition formation and improves overall performance.
After selecting a task location, a motion planner computes the shortest collision-free path to that location while avoiding other robots, obstacles, and other tasks. Although various approaches could be deployed, we adopt the A* search algorithm for simplicity. The path is dynamically recomputed whenever a potential collision is detected. Robots follow these updated paths to their assigned task locations. To improve coalition formation for higher-level tasks, the policy $\pi_{\phi}$ assigns actions by considering the long-term task execution plans of neighboring robots. Algorithm 1 outlines the motion control process for each robot guided by the learned policy $\pi_{\phi}$.
@@ 7      $(\bar{p}_{1},\cdots,\bar{p}_{n})$ = MotionPlanning ($p_{i,t},a_{i,t}$)
We also employ an intention map [23] that encodes each robot’s task location and collision-free path. This intention map is shared among neighboring robots to facilitate coalition formation. Inspired by the encoding scheme in [23], we use an exponentially decaying path representation, where the $i$-th position on the path is weighted by $e^{-\alpha i}$, with $\alpha\geq 0$, and embedded in the intention map.
IV-C Learning Task Allocation and Revision
To address this challenge, we enhance the MAPPO framework to train a policy capable of performing both task allocation and revision. The policy is trained to evaluate whether a robot’s current task allocation requires adjustment based on evolving local observations and information shared by neighboring robots. As demonstrated in §V, this integrated approach enables our framework to concurrently learn task allocation and revision, ensuring optimization of the overall objective in the Dec-POMDP environment.
Unlike existing approaches [23, 22], where task revisions occur only after a robot reaches its assigned task location and either completes it or fails due to the lack of a coalition, our trained policy proactively reallocates tasks whenever appropriate. In §V, we evaluate the performance improvements achieved by incorporating this task revision scheme compared to scenarios where such a scheme is absent.
V Evaluation
We conducted a series of simulations to evaluate (i) the performance of our proposed framework compared to existing methods, including the state-of-the-art market-based approach Project Manager-oriented Coalition Formation Algorithm (PCFA), [^2] The original PCFA [8] was modified to account for robots with limited view and communication ranges and to enable task reassignment. Details on these modifications are provided in the Appendix. (ii) its scalability with respect to the number of robots, and (iii) its generalizability across a range of diverse environments. Furthermore, we present an ablation study to underscore the critical role of task revision in dynamic coalition formation.
V-A Simulation Setup and Performance Metric
(a) Non-homogeneous tasks.
(b) Homogeneous tasks.
To evaluate our proposed method, we developed GyMRTA [^3] Our code will be made available soon., a Gym environment for discrete-space, discrete-time MRTA, and extended the existing MAPPO implementation from the EPyMARL [36] package. Our extensions include features such as task revision, action filtering, multi-dimensional observation and action spaces, support for custom state space, and tailored actor and critic Neural Network (NN) architectures. The hyperparameters of MAPPO, presented in Table III(a), remain constant across all simulations. Each episode is limited to 100 time steps ($T=99$) to ensure a finite episodic reward.
At the start of each episode, the initial positions of robots and tasks are randomly sampled. The environment is categorized based on task distribution as either homogeneous or non-homogeneous. In the homogeneous scenario, tasks are evenly distributed, maintaining a constant density across the grid. Conversely, the non-homogeneous scenario features uneven task distribution, resulting in variations in task density. Table II details the initial number of tasks per level across multiple task settings used in simulations.
TABLE II: Task
settings.
Setting Name Tasks
1 2 3
Easy
E1 15 - -
E2 10 - -
E3 5 5 -
Medium
M1 6 4 4
M2 4 3 3
M3 2 2 1
Hard
H1 - 5 5
H2 - - 10
H3 - - 5
Homogeneous
4xM2 16 12 12
In our simulations, we examined both homogeneous and non-homogeneous task distributions. Among the various possible non-homogeneous configurations, we selected the random corner patch distribution: the grid is divided into four equal-size corner patches, and at the start of each episode, one patch is randomly chosen to contain all tasks. Within the chosen patch, task density remains constant, while it is zero elsewhere. This setup evaluates whether the policy enables robots to cooperate in task execution while effectively locating tasks.
V-B Results
TABLE III:

### v2

@@ Abstract
We propose a decentralized, learning-based framework for dynamic coalition formation in Multi-Robot Task Allocation (MRTA). Our approach extends Multi-Agent Proximal Policy Optimization (MAPPO) by integrating spatial action maps, robot motion planning, intention sharing, and task allocation revision to enable effective and adaptive coalition formation. Extensive simulation studies confirm the effectiveness of our model, enabling each robot to rely solely on local information to learn timely revisions of task selections and form coalitions with other robots to complete collaborative tasks. Additionally, our model significantly outperforms existing methods, including a market-based baseline. Furthermore, we evaluate the scalability and generalizability of the proposed framework, highlighting its ability to handle large robot populations and adapt to scenarios featuring diverse task sets.
@@ I Introduction
Coalition formation, widely observed in nature, such as in ant colonies and beehives, inspires researchers and engineers to design simple robots capable of strategic cooperation. In multi-robot systems, coalitions emerge when the collective actions of robots exceed what they could accomplish individually, as illustrated in Fig. 1. In this work, we focus on developing policies for a team of robots capable of performing tasks that require coalition in dynamic environments. In such scenarios, robots must learn to coordinate by leveraging local information gathered from their own sensors and shared data from neighboring robots to effectively execute multi-robot tasks.
Multi-Robot Task Allocation (MRTA) is a research theme directly relevant to our work, focusing on the challenge of assigning and executing multiple tasks across a team of robots to optimize a collective objective over an extended time horizon. We focus on MRTA problems where: (i) each robot is capable of executing only one task at a time, (ii) tasks may require multiple robots to collaborate simultaneously for completion, (iii) robots must be adjacent to a task to execute it, necessitating movement to the task location beforehand, and (iv) completed tasks are removed, and new tasks continuously spawn in the environment. This problem formulation falls under the Single-Task robots, Multi-Robot tasks, Time-extended Assignment (ST-MR-TA) category of MRTA, as defined by the taxonomy in [1]. Relevant applications include transportation [2] and service dispatch scenarios [3, 4] such as firefighting and large-scale disaster response.
Figure 1: The firetruck dispatch as an illustrative example of a dynamic coalition formation problem: a team of firetrucks must form coalitions to tackle tasks, such as extinguishing fires, that cannot be accomplished individually.
@@
We propose a learning-based framework for multi-robot dynamic coalition formation to address problems within the ST-MR-TA MRTA category. In this framework, each robot utilizes a receding horizon-type planning strategy to generate a long-term plan for task execution, which is repeatedly revised as necessary in subsequent time steps. Each robot shares its plan with other robots that are within its communication range. This combination of task reassessment and sharing of the task execution plan enables dynamic coalition formation. Our contributions are summarized as follows:
@@
    Each robot’s policy is modeled as an end-to-end convolutional neural network based on the U-Net architecture [9]. The network processes information on nearby tasks and other robots within the robot’s sensing range and integrates the intentions of neighboring robots. It is trained using the MAPPO algorithm with the parameter-sharing technique.
@@
    Through extensive simulations, we demonstrate that our approach outperforms existing methods including a market-based approach, Project Manager-oriented Coalition Formation Algorithm (PCFA). Additionally, we demonstrate the scalability of models up to 1000 robots and their generalizability across a wide range of task settings.
@@ II Related Work
Solutions to ST-MR-TA MRTA include optimization-based [10, 11, 12, 13], trait-based [14, 15], and market-based [6, 8, 16, 7] methods. Among decentralized solutions for MRTA, market-based approaches are particularly popular. Specifically, PCFA [8] is a market-based solution designed for coalition formation in homogeneous teams with dynamic tasks. Due to the similarity in problem formulation, we selected PCFA as the baseline for our simulations reported in §V.
Reinforcement Learning (RL) focuses on learning a policy (i.e., a mapping from observations to actions) that maximizes a robot’s total reward in a Markov decision process. It has been widely applied to various robotics problems, particularly in object manipulation [17, 18, 19, 20, 21]. While many approaches aim to directly control robot actuators, studies such as [17, 18, 22] propose the use of an abstract action map. In this framework, the policy selects a task at each time step, while a motion planner handles the low-level actuator control to navigate the robot to the task location. This abstraction alleviates the learned policy from handling low-level control, allowing it to concentrate on long-term planning.
@@
Several studies have applied MARL to MRTA [29, 34, 35, 23]. For example, [29, 34] specifically address MRTA problems involving multi-robot tasks. The MRTA problem was introduced to the MARL community as a benchmark in [29], where the authors released a Python package called Level-Based Foraging (LBF) [29]. In LBF, each object is assigned a level, and it can only be picked up by a group of robots if the sum of the robots’ levels meets or exceeds the object’s level. Although LBF has become a popular MARL benchmark [36, 37, 38], it lacks support for dynamic object spawning and abstract action spaces (e.g., selecting task positions on the grid). In [36], several MARL algorithms were evaluated across popular benchmarks, including LBF (in fully observable scenarios only), and the EPyMARL package was released for community use. We expand their implementation by incorporating robot communication and additional features, which are detailed further in §IV.
@@
Figure 2: MRTA Environment-Robot Interaction: On the right, a depiction of the environment is shown along with the state representation $s_{t}$ used by the critic during training. Each robot receives an observation $o_{i,t}$ consisting of multiple channels. The observation is then fed into the policy $\pi_{\phi}$ and filtered to consider only visible tasks. A task is selected from the policy output, and the motion planner finds a feasible path to it.
@@ III Problem Description
Each task is associated with a positive integer $l_{i}\in\{1,\dots,L_{max}\}$, referred to as the task’s level, which represents the number of robots required to execute it. Here, $L_{max}$ denotes the maximum task level. We consider a dynamic task set: tasks are removed from the environment once completed by robots, and new tasks can appear at any time step in unoccupied cells. At time step $t$, the set of available tasks is represented as $\mathbb{T}_{t}=\{(q_{i},l_{i})\}_{i=1}^{T_{t}}$, where $q_{i}$ denotes the location of the $i$-th task, $l_{i}$ indicates its level, and $T_{t}$ represents the total number of tasks available.
@@
To design a decentralized policy under partial observability, we model the problem as a Decentralized Partially-Observable Markov Decision Process (Dec-POMDP) [39, 40]. A Dec-POMDP is defined by the tuple $(N,\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},O,R)$, where $N$ is the number of robots, $\mathcal{S}$ represents the state space of the environment, and $\mathcal{A}$ and $\mathcal{O}$ denote the action and observation spaces of individual robots, respectively. The state transition function $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}_{+}$ defines the probability $\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$ of transitioning to the next state $s_{t+1}\in\mathcal{S}$, given the current state $s_{t}\in\mathcal{S}$ and the joint action $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{N,t})\in\mathcal{A}^{N}$ of all robots. The observation function $O:\mathcal{S}\times\{1,\dots,N\}\to\mathcal{O}$ provides each robot $i$ with an observation $o_{i,t}=O(s_{t},i)$, mapping the state space $\mathcal{S}$ and robot index $i$ to the observation space $\mathcal{O}$. Lastly, the shared reward function is represented by $R:\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}$.
@@ IV Task Allocation Policy for Dynamic Coalition Formation
We elaborate on the two main contributions of this work: (i) the design of decentralized task allocation policy structure, and (ii) the development of a learning-based algorithm within the Dec-POMDP framework to identify the optimal policy. Our proposed policies are designed to select tasks for execution, assess whether a robot’s current task allocation requires revision, and determine new task assignments based on the robot’s evolving local information. Once a task is assigned, each robot performs low-level control to move to a neighboring cell along a collision-free path within a single time step. In the subsequent time step, the policy reassesses the task allocation based on updated local information, redirecting the robot to a new task if necessary. This iterative process ensures that the overall collective objective is optimized.
IV-A Dec-POMDP and Task Allocation Policy
IV-A1 Reward Function
@@
Higher-level tasks, which require coalitions of multiple robots, are more challenging to complete. To prevent robots from favoring lower-level tasks, we prioritize higher-level tasks by rewarding robots with a value equal to the square of the task level. Specifically, the reward at each time step $t$ is defined as $r_{t}=\sum_{(q_{i},l_{i})\in\mathbb{T}_{t}\backslash\mathbb{T}_{t+1}}l_{i}^{2}$, where $\mathbb{T}_{t}\backslash\mathbb{T}_{t+1}$ represents the set of tasks completed at time step $t$, and $l_{i}$ denotes the level of task $i$ [^1] Once the coalition requirement is met, we assume that the robots can complete each task within a single time step.
IV-A2 State and Observation Spaces
@@
We adopt a 2D occupancy grid representation segmented into multiple channels (see Table I), enabling the use of Convolutional Neural Network (CNN) architectures for policy design, similar to approaches in [35, 17, 23]. This differs from the representation used in LBF [29], which encodes the state as a vector of robots and their positions. In our framework, the state is defined as multiple separate channels as $s_{t}\triangleq(\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t},\mathcal{C}^{\text{intent}}_{t})$, where $\mathcal{C}^{\text{intent}}_{t}=\mathcal{C}^{\text{intent}}_{1,t}\times\cdots\times\mathcal{C}^{\text{intent}}_{N,t}$. [^2] To simplify the design and improve the efficiency of critic model training, our implementation consolidates $\mathcal{C}^{\text{intent}}_{t}$ into a single aggregated channel that combines the intention channels of all robots. Empirical studies indicate that this aggregation does not impact the training outcomes.
@@
The observation $O(s_{t},i)$ for robot $i$ is derived by projecting and cropping $s_{t}$ onto areas centered around the robot’s position, determined by its view range $R_{\text{view}}$ and communication range $R_{\text{comm}}$. Specifically, the robot intention channel $\mathcal{C}^{\text{intent}}_{t}$ includes $\mathcal{C}^{\text{intent}}_{j,t}$ from robot $j$ if $j$ is within $R_{\text{comm}}$ cells of robot $i$. Meanwhile, the other channels, $\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t}$, are obtained by cropping $s_{t}$ according to $R_{\text{view}}$.
TABLE I: Description of CNN input channels.
@@
IV-A3 Action Space, Motion Planner, and Intention Map
Following a similar representation as in [17, 18], a robot’s action – determined by the task allocation policy $\pi_{\phi}$ – is defined as the location of its next task. Robots can select task locations only if they are within its view range; if no tasks are within view range, all grid locations within communication range are available for action selection. This constraint provides prior information, reducing training time while encouraging robots to commit to specific tasks.
After selecting a task location, a motion planner computes the shortest collision-free path to that location while avoiding other robots, obstacles, and other tasks. We adopt the A* search algorithm for simplicity. The path is dynamically recomputed whenever a potential collision is detected. Robots follow these updated paths to their assigned task locations. To improve coalition formation for higher-level tasks, the policy $\pi_{\phi}$ assigns actions by considering the long-term task execution plans of neighboring robots.
@@ 7      $(\bar{p}_{1},\cdots,\bar{p}_{n})$ = MotionPlanning ($p_{i,t},a_{i,t}$)
8      Generate an intention map with $(\bar{p}_{1},\cdots,\bar{p}_{n})$ and share it with neighboring robots.
@@
We also employ an intention map [23] that encodes each robot’s task location and collision-free path. Inspired by the encoding scheme in [23], we use an exponentially decaying path representation, where the $i$-th position on the path is weighted by $\alpha^{-i}$, with $\alpha=2/3$ in our simulation, and embedded in the intention map. This intention map is shared among neighboring robots to facilitate coalition formation. Algorithm 1 outlines the motion control process for each robot guided by the learned policy $\pi_{\phi}$.
IV-B Optimal Task Allocation Policy
IV-B1 Multi-Agent Proximal Policy Optimization (MAPPO)
To compute the optimal policy, we adopt MAPPO [27], a CTDE algorithm designed for MARL. MAPPO enables robots to share experiences during the learning phase, which is particularly beneficial for training policies in teams of homogeneous robots. MAPPO is an Actor-Critic algorithm comprising two distinct networks: the actor $\pi_{\phi}$ and the critic $V_{\theta}$. The actor outputs a distribution over actions based on each robot’s observations, formally defined by the policy $\pi_{\phi}:\mathcal{O}\times\mathcal{A}\to\mathbb{R}_{+}$, where $\phi$ represents the parameters of the policy network. These parameters $\phi$ are shared across all robots, a technique known as parameter sharing [41], which improves training efficiency, especially in cooperative scenarios involving homogeneous robots. The critic computes an approximate value function $V_{\theta}:\mathcal{S}\to\mathbb{R}$, parameterized by $\theta$, to estimate the expected discounted sum of future rewards $\mathbb{E}\big{[}{\sum_{\tau=t}^{T}\gamma^{\tau-t}R(s_{\tau},\mathbf{a}_{\tau})}\big{]}$.
The critic $V_{\theta}$ is trained to accurately estimate future rewards, while the actor $\pi_{\phi}$ is trained to select actions that maximize the critic’s estimate. In our framework, both the actor $\pi_{\phi}$ and the critic $V_{\theta}$ are implemented using U-Nets [9], a CNN-based architecture. For this implementation, the actor’s input (observations) and output (action distribution) as well as the critic’s input (state), are represented as images (3D arrays). The U-Net architecture is particularly advantageous due to its ability to capture both local spatial relationships and global context within the input image, making it well-suited for policy learning.
TABLE II: